# A gluing construction of $D_{k}$ ALF gravitational instantons and existence of non-holomorphic minimal spheres

### v1

@@ Abstract.
This note extends the construction of $D_{k}$ ALF gravitational instantons in Singer–Schroers [SS21] to a new case where the nonlinear superposition is given by the $D_{1}$ Atiyah–Hitchin metric and $k-1$ copies of $A_{0}$ Taub-NUT metrics. We then give a general class of ALF spaces such that each of them contains a non-holomorphic minimal sphere. Together with Foscolo’s construction [Fos19] this gives a large class of $K3$ surfaces containing non-holomorphic minimal spheres.
@@ 1. Introduction
A gravitational instanton is a complete noncompact 4-dimensional hyperKähler manifold with $L^{2}$ curvature decay at infinity. There has been a long history in the study of gravitational instantons and a lot of recent developments have been made. Gravitational instantons arise as building blocks for degeneration of K3 surfaces [GW00, Fos19, HSVZ22, CVZ20], which makes them powerful tools in probing $K3$ surfaces near degeneration limits.
The gravitational instantons have been classified into six types, ALE / F / G / H / G* / H*, based on the end structure at infinity [Min10, CC21, SZ22]. In this paper we will discuss the ALF gravitational instantons, which is modeled at infinity by quotients of circle bundles over $\mathbb{R}^{3}$ and has cubic volume growth. There are two families of ALF spaces depending on the monodromy at infinity being cyclic or dihedral, $A_{k}$ and $D_{k}$, labelled by non-negative integers $k$. In this paper we will discuss the construction of $D_{k}$ type ALF spaces via a nonlinear superposition of $D_{1}$ and $k-1$ copies of $A_{0}$ ALF spaces, and using this we will show the existence of non-holomorphic minimal spheres in these spaces as well as in $K3$ surfaces.
This gluing construction extends the work by Schroers and Singer [SS21], who obtained $D_{k}$ ALF gravitational instantons by gluing the moduli space ${\mathcal{M}}_{2}^{0}$ of centered SU(2) monopoles (which is the only $D_{0}$ ALF space) with $k$ copies of Taub-NUT space (each of which is an $A_{0}$ ALF space). As commented in their paper, this construction can be extended to other cases. In this paper we work on the case where the center piece is replaced by the Atiyah–Hitchin metric AH [AH88], which is defined to be the double cover of ${\mathcal{M}}_{2}^{0}$ and is a rotationally invariant $D_{1}$ space. A rough statement of the theorem is given below and a more explicit description will be given at Theorem 3.
@@
For any $k\geq 1$ and any set $\{0,p_{1},\dots,p_{k}\}\in\mathbb{R}^{3}$ of $k+1$ distinct points, there exists $\epsilon_{0}>0$, such that for any $\epsilon\in(0,\epsilon_{0})$ there exists a $D_{k+1}$ ALF metric $g_{\epsilon}$ which is a nonlinear superposition of the Atiyah–Hitchin metric centered at 0 and $k$ copies of Taub-NUT metrics centered at $p_{i}/\epsilon,1\leq i\leq k$.
There has been a lot of progress in the analysis and construction of ALF spaces [Min09, Min10, CC19, BM11]. The $A_{k}$ type ALF has been well understood as a multi-Taub-NUT space via gluing construction [Min11, GH78]. And $D_{k}$ spaces have been constructed via twistor methods [Hit84] and desingularization [BM11, Auv18, Dan93]. The construction in [SS21] gives an explicit gluing constructions of $D_{k}$ type ALF based on Sen’s proposal [Sen97], and the construction follows the physical interpretation of modeling electronically charged particles like protons and electrons [AMS12]. There are also works by Cherkis, Kapustin, Hitchin, Roček et.al [CH05, CK98, CK99, LRc88, IRc96] describing ALF spaces as moduli space of monopoles with prescribed singularities. The construction in [SS21] gave $D_{k}$ ALF when the $k$ Taub-NUT singularities are well separated from the origin, while the ones constructed in this paper can be viewed as the limit of these $D_{k}$ spaces when one Taub-NUT singularity approaches zero.
It can be potentially further generalized, as suggested in [SS21], to cases of composing a $D_{j}$ ALF space with various $A_{\nu_{i}-1}$ ALF to generate other $D_{k^{\prime}}$ spaces, where $k^{\prime}=j+\sum\nu_{i}$, which would give other limiting cases when more TN singularities travel to zero or to each other. Since in this paper we are mainly interested in using the property of Atiyah–Hitchin space, we will only work on one generalization where the center piece is exactly given by AH.
Our construction follows the same procedure and the analysis is done following [SS21], as the change of the center piece from ${\mathcal{M}}_{0}^{2}$ to $AH$ does not change the asymptotic behavior. The mapping properties of such metrics can be analyzed using the fiber boundary calculus, or $\phi-$calculus [MM98, Vai01, HHM04], which captures the asymptotic metric structure of ALF spaces.
As a consequence of the construction, we obtain a family of $D_{k}$ spaces that, with properly chosen marked points, degenerate to the Atiyah–Hitchin space as the Gromov Hausdorff limit. One special feature of the AH space is that it contains a stably minimal sphere that is not holomorphic with any of the complex structures in the hyperKähler rotation [MW93]. With this construction we can obtain such special sphere in all nearby $D_{k}$ space.
@@
For any $k>1$, there is a family of $D_{k}$ ALF metrics $\{g_{\epsilon}\}_{\epsilon\in(0,\epsilon)}$, such that each of them contains a minimal sphere that is not holomorphic with respect to any of the compatible complex structure.
In [Fos19] Foscolo constructed a collapsing sequence of K3 surfaces using ALF spaces as part of building blocks. As a corollary he used Atiyah–Hitchin space as the building block to show the existence of minimal non-holomorphic sphere in some K3 surface near the degeneration limit. Now with more ALF spaces constructed above with non-holomorphic spheres, we extend this result of Foscolo and show that there is a neighborhood in the moduli space of K3 surfaces such that each of these has a special sphere.
@@
There is an open neighborhood in the moduli space of hyperKähler metrics on $K3$, such that each metric in this neighborhood contains a stably minimal sphere that is not holomorphic with respect to any compatible complex structures.
This corollary is connected to one interpretation of the Weak Gravity Conjecture [AHMNV07], which predicts the existence of “non-BPS” particles. In mathematics such particles can be interpreted as non-holomorphic submanifolds in Calabi–Yau manifolds in the large complex limit [HLM+20, DLMS20, LSVY23]. The corollary above can be seen as one of the examples.
The paper is organized in the following order: Section 2 describes the two building blocks, the multi-Taub-NUT space and the Atiyah-Hitchin metric, then gives the gluing construction and perturbative argument to find the hyperKähler metric. Section 3 proves the existence of non-holomorphic spheres in some $D_{k}$ ALF spaces and $K3$ surfaces.
Acknowledgement: The author is partially supported by NSF DMS-2305363
@@ 2. The gluing construction of $D_{k}$ ALF spaces
The gluing construction is a modification from [SS21], and we describe the construction below. The basic idea is to first construct an adiabatic Gibbons–Hawking Ansatz with symmetric Taub–NUT singularities which give the correct $D_{k}$ asymptotic topology at infinity after the quotient of an involution action. After that we glue in a double branched cover of Atiyah Hitchin metric in the center. This produces an approximate hyperKähler metric after involution. Finally we solve for the actual hyperKähler metric by perturbation of hyperKähler triples. The set up is similar to [SS21] where the center piece is a double branched cover of ${\mathcal{M}}_{2}^{0}$. We also need to modify the coefficients in the adiabatic Gibbons–Hawking ansatz to be compatible with $AH$ metric. On the other hand, the asymptotic behavior of ${\mathcal{M}}_{2}^{0}$ and Atiyah Hitchin metric are similar since $AH$ is the double cover of ${\mathcal{M}}_{2}^{0}$, the analysis used for perturbing the approximate triple to get the exact solution can be applied directly.
Now we describe the two pieces needed for gluing
@@ 2.1. The multi-Taub-NUT space $M_{\epsilon}$
Fix $\delta\in(0,1/2)$, and let $\epsilon\in(0,\delta^{2})$ which will go to 0 in the limit later. Fix $p_{1},\dots,p_{k}\in\mathbb{R}^{3}$ sufficiently large, we consider a family of 4-manifolds $M_{\epsilon}$, each of which is a singular circle fibration over $\{x\in\mathbb{R}^{3}:|x|>2\epsilon\}$. We consider the following adiabatic family of Gibbon–Hawking ansatz on $M_{\epsilon}$
@@
where $\alpha$ is a connection 1-form of the circle bundle, and $*_{\epsilon}$ is the Hodge star with respect to the flat metric $\frac{|dx|^{2}}{\epsilon^{2}}$ on the base |$\mathbb{R}^{3}_{|x|>2\epsilon}$; the singular harmonic function $h_{\epsilon}$ on |$\mathbb{R}^{3}_{|x|>2\epsilon}$ is given by
(2) $h_{\epsilon}=1-\frac{\epsilon}{|x|}+\sum_{v=1}^{k}\left(\frac{\epsilon}{2|x+p_{v}|}+\frac{\epsilon}{2|x-p_{v}|}\right).$
We take the singular fibration associated with the Gibbons–Hawking ansatz (1) by
@@
and note the fibration is only singular at $P:=\{\pm p_{1},\dots,\pm p_{2}\}$. This ansatz is modified from [SS21, equation (1.4)] by replacing the coefficients of $\frac{1}{|x|}$ from $2\epsilon$ to $\epsilon$ which will match the the Atiyah–Hitchin metric in the center later. At ±$\pm p_{i}$ this metric has a Taub-NUT type singularity, which is natural since $A_{k}$ ALF spaces can be viewed as a multi-Taub-NUT space.
@@
(4) $h_{\epsilon}=1+\frac{\epsilon(2k-2)}{2|x|}+O(\epsilon|x|^{-3}),\ |x|\gg 1$
which, after quotient by the involution $\iota$, gives the asymptotic topology of a dihedral $D_{k+1}$ ALF.
We recall one fact about the behavior of $g_{\epsilon}$ when lifted to the compactified space, where the detailed proof was given in [SS21, Proposition 3.4].
@@
The fibration $\phi:M_{\epsilon}\rightarrow\mathbb{R}^{3}_{|x|>2\epsilon}$ can be lifted to the radially-compactified base $\overline{\mathbb{R}^{3}}\setminus\{|x|\leq 2\epsilon\}$. The radial compactification of $\mathbb{R}^{3}$ is obtained by adding a sphere $\mathbb{S}^{2}=\partial\overline{\mathbb{R}^{3}}$ at infinity. We use the inverted polar coordinates $\rho=1/|x|,y=\rho x$ near the boundary, and the metric $g_{\epsilon}$ in (1) lifts to a fiber boundary metric of the form
@@
where $h_{\epsilon}$ and $\bar{\alpha}=*_{\epsilon}dh_{\epsilon}$ lift to be smooth on ¯$\bar{\mathbb{R}}^{3}\setminus\{|x|\leq 2\epsilon\}$ (one can check this by changing variables in (2)).
@@ 2.2. The branched cover $\widehat{AH}_{\epsilon}$
[…]
@@
where $x^{\prime}$ is bijectively related to $\tilde{x}$ in an implicit way. In particular, $AH$ and its branched cover $\widehat{AH}$ are both strongly ALF spaces by [SS21, Definition 3.5] and the asymptotic analysis at infinity can be applied the same way.
To connect with the adiabatic Gibbons–Hawking ansatz constructed before, we consider the following metric
@@
$h_{\epsilon}^{\prime}=1-\frac{1}{|x^{\prime}|}+\sum_{v=1}^{k}\frac{\epsilon}{|p_{v}|},\ \alpha^{\prime}=*dh_{\epsilon}^{\prime}.$
We call the associated circle fibration total space $\widehat{AH}_{\epsilon}$ and the fibration

### v2

@@ Abstract.
This note extends the construction of $D_{k}$ ALF gravitational instantons in Schroers–Singer [SS21] to a new case where the nonlinear superposition is given by the $D_{1}$ Atiyah–Hitchin metric and k1k-1 copies of $A_{0}$ Taub-NUT metrics. We then give a general class of ALF spaces such that each of them contains a non-holomorphic minimal sphere. Together with Foscolo’s construction [Fos19] this gives a large class of $K3$ surfaces containing non-holomorphic minimal spheres.
@@ 1. Introduction
A gravitational instanton  [^1] The term ”gravitational instanton” is debated and might mean a different object for other authors. In this paper we only discuss the hyperKähler case. is a complete noncompact real 4-dimension hyperKähler manifold with $L^{2}$ curvature decay at infinity. There has been a long history in the study of gravitational instantons and a lot of recent developments have been made. Gravitational instantons arise as building blocks for degeneration of K3 surfaces [GW00, Fos19, HSVZ22, CVZ20], which makes them powerful tools in probing $K3$ surfaces near degeneration limits.
The gravitational instantons have been classified into six types, ALE/ ALF/ALG/ALH/ALG*/ALH*, based on the end structure at infinity [Min10, CC21, SZ24]. In this paper we will discuss the ALF gravitational instantons, which is modeled at infinity by quotients of circle bundles over $\mathbb{R}^{3}$ and has cubic volume growth. There are two families of ALF spaces depending on the monodromy at infinity being cyclic or dihedral, $A_{k}$ and $D_{k}$, labelled by non-negative integers $k$. In this paper we will discuss the construction of $D_{k}$ type ALF spaces via a nonlinear superposition of $D_{1}$ and k1k-1 copies of $A_{0}$ ALF spaces, and using this we will show the existence of non-holomorphic minimal spheres in these spaces as well as in $K3$ surfaces.
This gluing construction extends the work by Schroers and Singer [SS21], where the authors constructed $D_{k}$ ALF gravitational instantons by gluing the moduli space ${\mathcal{M}}_{2}^{0}$ of centered SU(2) monopoles (which is the only $D_{0}$ ALF space) with $k$ copies of Taub-NUT space (each of which is an $A_{0}$ ALF space). As commented in their paper, this construction can be extended to other cases. In this paper we work on the case where ${\mathcal{M}}_{2}^{0}$ is replaced by the Atiyah–Hitchin metric AH [AH88], which is defined to be the double cover of ${\mathcal{M}}_{2}^{0}$ and is a rotationally invariant $D_{1}$ space. Gluing the AH metric with k1k-1 copies of Taub-NUT space (one copy fewer than [SS21]), we get a $D_{k}$ space. A rough statement of the theorem is given below and a more explicit description will be given at Theorem 3.
@@
Given any $k\geq 2$ and any set of $k$ distinct points in $\mathbb{R}^{3}$ denoted as $\{0,p_{1},\dots,p_{k-1}\}$ , there exists $\epsilon_{0}>0$, such that for any $\epsilon\in(0,\epsilon_{0})$ there exists a $D_{k}$ ALF metric $g_{\epsilon}$ which is a nonlinear superposition of the Atiyah–Hitchin (AH) metric centered at 0 and k1k-1 copies of Taub-NUT metrics (TN) centered at $p_{i}/\epsilon,1\leq i\leq k-1$.
There has been a lot of progress in the analysis and construction of ALF spaces [Min09, Min10, CC19, BM11]. The $A_{k}$ type ALF has been well understood as a multi-Taub-NUT space via gluing construction [Min11, GH78]. And $D_{k}$ spaces have been constructed via twistor methods [Hit84] and desingularization [BM11, Auv18, Dan93]. The construction in [SS21] gives an explicit gluing constructions of $D_{k}$ type ALF based on Sen’s proposal [Sen97], and the construction follows the physical interpretation of modeling electrically charged particles like protons and electrons [AMS12]. There are also works by Cherkis, Kapustin, Hitchin, Roček et.al [CH05, CK98, CK99, LR88, IR96] describing ALF spaces as moduli space of monopoles with prescribed singularities. The construction in [SS21] gave $D_{k}$ ALF when the $k$ Taub-NUT singularities are well separated from the origin, while the ones constructed in this paper can be viewed as the limit of these $D_{k}$ spaces when one of these Taub-NUT singularities approaches zero.
It can be potentially further generalized, as suggested in [SS21], to cases of composing a $D_{j}$ ALF space with various $A_{\nu_{i}-1}$ ALF to generate other $D_{k^{\prime}}$ spaces, where $k^{\prime}=j+\sum\nu_{i}$, which would give other limiting cases when more TN singularities travel to zero or to each other. Since in this paper we are mainly interested in using the property of Atiyah–Hitchin space, we will only work on the case where the center piece is exactly given by AH.
Our construction follows the same procedure and the analysis is done following [SS21], as the change of the center piece from ${\mathcal{M}}_{0}^{2}$ to $AH$ does not change the asymptotic behavior. The mapping properties of such metrics can be analyzed using the fiber boundary calculus, or $\phi$-calculus [MM98a, Vai01, HHM04], which captures the asymptotic metric structure of ALF spaces.
As a consequence of the construction, we obtain a family of $D_{k}$ spaces that, with properly chosen marked points, degenerate to the Atiyah–Hitchin space as the Gromov Hausdorff limit. One special feature of the AH space is that it contains a minimal sphere that is not holomorphic with any of the complex structures in the hyperKähler rotation [MW93]. In the same paper it was shown that this minimal sphere is stable, i.e. the second variation of the area functional is nonnegative. In fact this sphere is shown to be ”strongly stable” and area minimizing by Tsai and Wang [TW22]. This special sphere was also recently discussed in the compactness of Fueter sections [EL24]. With our gluing construction we can obtain such special sphere in all nearby $D_{k}$ space.
@@
For any $k\geq 2$, there is a family of $D_{k}$ ALF metrics $\{g_{\epsilon}\}_{\epsilon\in(0,\epsilon)}$, such that each of them contains a stable minimal sphere that is not holomorphic with respect to any of the compatible complex structures in the hyperKähler rotation.
There has been a lot of recent developments relating to minimal surfaces in the context of gravitational instantons. Lotay and Oliveira [LO24] classified all $\mathbb{S}^{1}$-invariant minimal surfaces in Gibbons–Hawking ansatz. Foscolo and Trinca [FT24] recently constructed new non-holomorphic minimal 2-spheres in multi-Taub–NUT spaces and K3 surfaces using Scherk surface and holomorphic cigars, yet those 2-spheres are unstable. For more general discussion regarding (non-)holomorphic minimal surfaces, we also refer to [Are98, Don06, MW06].
In [Fos19] Foscolo constructed a collapsing sequence of K3 surfaces using ALF spaces as part of the building blocks. As a corollary he used Atiyah–Hitchin space as the building block to show the existence of minimal non-holomorphic sphere in some K3 surface near the degeneration limit. Now with more ALF spaces constructed above with non-holomorphic spheres, we extend this result of Foscolo and show that there is a neighborhood in the moduli space of K3 surfaces such that each of these has a special sphere.
@@
There is an open neighborhood in the moduli space of hyperKähler metrics on $K3$, such that each metric in this neighborhood contains a stable minimal sphere that is not holomorphic with respect to any compatible complex structures.
This corollary is connected to one interpretation of the Weak Gravity Conjecture [AHMNV07], which predicts the existence of “non-BPS” particles. In mathematics such particles can be interpreted as non-holomorphic submanifolds in Calabi–Yau manifolds in the large complex structure limit [HLM+20, DLMS20, LSVY23]. The corollary above can be seen as one of the examples.
The paper is organized in the following order: Section 2 describes the two building blocks, the multi-Taub-NUT space and the Atiyah–Hitchin metric, then gives the gluing construction and perturbative argument to find the hyperKähler metric. Section 3 proves the existence of non-holomorphic spheres in some $D_{k}$ ALF spaces and $K3$ surfaces.
Acknowledgement: The author would like to thank Lorenzo Foscolo, Rafe Mazzeo, Michael Singer and Saman Habibi Esfahani for helpful discussions, and the anonymous referee for the careful reading and valuable feedback. The author is partially supported by NSF DMS-2305363.
@@ 2. The gluing construction of $D_{k}$ ALF spaces
The gluing construction is a modification from [SS21], and we describe the construction below. The basic idea is to first construct an adiabatic Gibbons–Hawking Ansatz which is a singular circle fibration over $\mathbb{R}^{3}$ outside a compact set and has 2(k1)2(k-1) $\mathbb{Z}_{2}$-symmetric Taub–NUT singularities $\{\pm p_{i}/\epsilon\}_{i=1}^{k-1}\subset\mathbb{R}^{3}$. Then we glue in a double branched cover of Atiyah–Hitchin metric, which is also a circle fibration, into the compact part of $\mathbb{R}^{3}$. After the quotient of the involution action $x\rightarrow-x$ in $\mathbb{R}^{3}$, this approximate metric gives the correct $D_{k}$ asymptotic topology at infinity (given by the multi-Taub–NUT metrics) and is the smooth Atiyah–Hitchin metric near 0. However the metric is only approximate hyperKähler. Finally we solve for the actual hyperKähler metric by perturbation of the hyperKähler triple.
The set up is similar to [SS21] where the center piece is a double branched cover of ${\mathcal{M}}_{2}^{0}$. We also modify the coefficients in the adiabatic Gibbons–Hawking ansatz so that it is compatible with the $AH$ metric. Except those changes, the asymptotic behavior of ${\mathcal{M}}_{2}^{0}$ and Atiyah–Hitchin metric are similar since $AH$ is the double cover of ${\mathcal{M}}_{2}^{0}$, therefore the analysis used for perturbing the approximate triple to get the exact solution can be applied directly.
Now we describe the two pieces needed for gluing.
@@ 2.1. The multi-Taub-NUT space $M_{\epsilon}$
Fix $\delta\in(0,1/2)$, and let $\epsilon\in(0,\delta^{2})$ be a small parameter which will go to 0 in the limit later. Fix a set of distinct points $p_{1},\dots,p_{k-1}\in\{x\in\mathbb{R}^{3}:|x|>2\epsilon\}$, we consider a family of 4-manifolds $M_{\epsilon}$, each of which is a singular circle fibration over $\{x\in\mathbb{R}^{3}:|x|>2\epsilon\}$. We consider the following adiabatic family of Gibbon–Hawking ansatz on $M_{\epsilon}$
@@
where $\alpha$ is a connection 1-form of the circle bundle, and $_{\epsilon}$ is the Hodge star with respect to the flat metric $\frac{|dx|^{2}}{\epsilon^{2}}$ on the base $\mathbb{R}^{3}_{|x|>2\epsilon}$; the singular harmonic function $h_{\epsilon}$ on $\mathbb{R}^{3}_{|x|>2\epsilon}$ is given by
(2) $h_{\epsilon}=1-\frac{\epsilon}{|x|}+\sum_{v=1}^{k-1}\left(\frac{\epsilon}{2|x+p_{v}|}+\frac{\epsilon}{2|x-p_{v}|}\right).$
We denote $\phi$ to be the singular fibration associated with the Gibbons–Hawking ansatz (1) by
@@
and note the fibration is only singular at $P:=\{0,\pm p_{1},\dots,\pm p_{k-1}\}$. This ansatz is modified from [SS21, equation (1.4)] by replacing the coefficients of $\frac{1}{|x|}$ from $2\epsilon$ to $\epsilon$ which will match the Atiyah–Hitchin metric in the compact part later. Note that, by a change of variables $\tilde{x}=x/\epsilon,q_{v}=p_{v}/\epsilon$, the harmonic function $h_{\epsilon}$ can be written as
$h_{\epsilon}=1-\frac{1}{|x|}+\sum_{v=1}^{k-1}\left(\frac{1}{2|x+q_{v}|}+\frac{1}{2|x-q_{v}|}\right)$
which gives a Gibbon–Hawking ansatz with Taub–NUT singularities at $q_{v}$, therefore this matches the name “multi-Taub–NUT space” for $A_{k-1}$ ALF spaces.
@@
(4) $h_{\epsilon}=1+\frac{\epsilon(2k-4)}{2|x|}+O(\epsilon|x|^{-3}),\ |x|\gg 1$
which, after quotient by the involution $\iota$, gives the asymptotic topology of a dihedral $D_{k}$ ALF.
Such behavior is characterized by “fiber boundary structures” [MM98b], and we recall the definition below. We will be working with a manifold $X$ with boundary $\partial X$, and a boundary defining function $\rho$ for $\partial X$ is a smooth function on $X$ such that $\rho|_{\partial X}=0$ and $d\rho$ is nowhere vanishing on $\partial X$.
@@
The fibration $\phi:M_{\epsilon}\rightarrow\mathbb{R}^{3}_{|x|>2\epsilon}$ can be lifted to the radially-compactified base $\overline{\mathbb{R}^{3}}\setminus\{|x|\leq 2\epsilon\}$. Here the radial compactification of $\mathbb{R}^{3}$ is obtained by adding a sphere $\mathbb{S}^{2}=\partial\overline{\mathbb{R}^{3}}$ at infinity. Using $x\in\mathbb{R}^{3}$ as the usual coordinates, the compactification introduces the inverted polar coordinates $\rho=1/|x|,y=\frac{x}{|x|}\in\mathbb{S}^{2}$ near the boundary, and the metric $g_{\epsilon}$ in (1) lifts to a fiber boundary metric of the form
@@
where $h_{\epsilon}$ and $\bar{\alpha}=*_{\epsilon}dh_{\epsilon}$ lift to be smooth on $\overline{\mathbb{R}^{3}}\setminus\{|x|\leq 2\epsilon\}$ (one can check this by changing variables in (2)). A detailed description was given in [SS21, Proposition 3.4].
@@ 2.2. The branched cover $\widehat{AH}_{\epsilon}$
We will glue the (double branched cover) of Atiyah–Hitchin metric into the center. Following the notation in [SS21], we denote by $\widehat{AH}$ the space $\mathbb{C}\mathbb{P}^{1}\times\mathbb{C}\mathbb{P}^{1}\setminus\text{anti-diagonal}$, or in coordinates:
@@
where $x^{\prime}$ is bijectively related to $\tilde{x}$ in an implicit way. In particular, $AH$ and its branched cover $\widehat{AH}$ are both strongly ALF spaces in the sense of [SS21, Definition 3.5]. Recall a $\phi$-metric $g$ is called “strongly ALF” if it is close to an exact Gibbons-Hawking ansatz of $A_{k}$ or $D_{k}$ ALF space $g_{ALF}$ modulo an error vanishing to infinite order, i.e. outside a compact set $|g-g_{ALF}|={\mathcal{O}}(\rho^{\infty})$ where $\rho$ is a boundary defining function for $g$ with respect to the $\phi$-structure. In particular, both $AH$ and $\widehat{AH}$ satisfy this condition, with an exponentially decaying error. Having this condition ensures the asymptotic analysis at infinity can be applied the same way. As pointed out in [SS21, Theorem C.4 and Remark C.5], when $g$ is strongly ALF, one can construct a “nicer” parametrix to solve the Poisson equation $\Delta_{g}u=f$. When the right hand side $f$ is smooth and decaying, one can get a smooth decaying solution $u$. On the other hand for a generic ALF metric one only expects polyhomogeneity rather than smoothness for $u$ in this case. This is used in the second step of the proof of Theorem 3, where the linearized operator is given by 12 copies of scalar Laplace operators, and one needs to iteratively solve away the error and make sure the solution is still smooth.
To connect the asymptotics of $g_{\widehat{AH}}$ with the adiabatic Gibbons–Hawking ansatz constructed in (1), we modify (7) to add the Taub-NUT singularities and get
@@
$h_{\epsilon}^{\prime}=1-\frac{1}{|x^{\prime}|}+\sum_{v=1}^{k-1}\frac{\epsilon}{|p_{v}|},\ \alpha^{\prime}=*dh_{\epsilon}^{\prime}.$
Here $|x^{\prime}|>\frac{1}{\delta}$ for $\delta\ll 1$. We call the associated circle fibration total space $\widehat{AH}_{\epsilon}$ and the fibration
@@
Later we will connect the large neck region $\{\frac{1}{\delta}<|x^{\prime}|<\frac{\delta}{\epsilon}\}$ of $\widehat{AH}_{\epsilon}$ to the region of $\{\frac{\epsilon}{\delta}<|x|<\delta\}$ in the multi-Taub-NUT $M_{\epsilon}$.